### v1

# On the Cut Locus of Submanifolds of a Finsler Manifold

Abstract.
In this paper, we investigate the cut locus of submanifolds in a Finsler manifolds, a natural generalization of Riemannian manifolds. We explore the deformation and characterization of the cut locus, extending the results of [BP21]. We also obtain a generalization of Klingenberg’s lemma for closed geodesic ([Kli59]) in the Finsler setting.
Key words and phrases:
cut locus, conjugate locus, closed geodesics, Finsler geometry, Finsler submanifolds, cone bundles
2020 Mathematics Subject Classification:
Primary: 53C22, 53B40; Secondary: 53C60
1. Introduction
In the field of Riemannian geometry, the study of geodesics plays a fundamental role in comprehending the geometric properties of the underlying space. Geodesics are (locally) the shortest paths between two points on a Riemannian manifold, akin to straight lines in Euclidean spaces. One important concept related to geodesics is the cut locus. The cut locus of a point (or a submanifold) of a Riemannian manifold is the set of points beyond which a distance minimizing geodesic from the point (or the submanifold) ceases to be distance minimal. Cut locus of a point, a notion introduced by Henri Poincaré [Poi05], has been extensively studied (see [Kob67] for a survey as well as [Buc77], [Mye35], [Wol79], and [Sak96]). Prior to Poincaré, its implicit appearance maybe traced back to a paper by Mangoldt [vM81]. Other articles such as [Whi35] and [Mye35, Mye36] describe topological properties of the cut locus. The exploration of the cut locus of a Riemannian manifold plays a crucial role in understanding the manifold’s underlying geometry, leading to many interesting results in Riemannian geometry. For instance, Rauch’s proof of the “Sphere theorem” [Rau59] heavily relies on estimates of distances to the cut locus. Additionally, it was realized that much of a manifold’s topological properties lie in its cut locus. A very good account of these results and methods are contained in articles [Kli59, Kob67, Wei68].
Finsler manifolds provide a generalization of Riemannian manifolds by considering a broader class of metrics that capture both the length and the direction of tangent vectors. The study of Finsler geometry has yielded significant insights and applications in various fields, including physics, optimization theory, and robotics. Essentially, a Finsler manifold is a manifold $M$ where each tangent space is equipped with a Minkowski norm, which is a norm not necessarily induced by an inner product. In sharp contrast to the Riemannian case, this norm induces canonical inner products that are not parameterized by points of $M$, but by directions in $TM$. Thus, one can think of a Finsler manifold as a space where the inner product does not only depend on where you are but also in which direction you are looking. While a Riemannian manifold is itself trivially a Finsler manifold, where the inner product at a point is the same in all the tangent directions, Finsler geometry represents quite a large step away from the Riemannian realm. Nonetheless, Finsler geometry contains analogs for many of the natural objects in Riemannian geometry. For example, the length of a curve, geodesics, curvature, connections, and covariant derivatives all generalize; however, normal coordinates do not [Run59]. In this article, we focus on investigating the cut locus of a submanifold within a Finsler manifold. For a Finsler manifold, the cut locus of a point has been extensively studied, see [War65, Wei68, Has73, ST16].
[…]
Theorem A (4.10).
[…]
  1. (1)
    $M\setminus\mathrm{Cu}(N)$ strongly deformation retracts onto $N$, and
  2. (2)
    $M\setminus N$ strongly deformation retracts onto $\mathrm{Cu}(N)$.
Along with the above result, we give two characterizations of the cut locus. The first one is in terms of the focal locus (4.6), and the second one is in terms of separating set (4.8). Recall that the focal locus is the collection of points where the normal exponential map fails to be of full rank (3.22), whereas the separating set of a submanifold is the collection of all points which can be joined by more than one minimizing geodesics from the submanifold (4.1). The second characterization is particularly important in calculations, as shown in [Pra23].
Theorem B (4.6).
[…]
  1. (1)
    There exists two distinct $N$-segments joining $N$ to $q$, i.e, $q\in\mathrm{Se}(N)$, or
  2. (2)
    $q$ is the first focal point of $N$ along $\gamma$.
Theorem C (4.8).
[…]
In Finsler manifolds, geometric objects possess a two-sided nature, exhibiting both forward and backward characteristics. This duality arises due to the inherent asymmetry of the distance function in Finsler geometry. Consequently, Klingenberg’s celebrated result on closed geodesics loops in Riemannian geometry ([Kli59]), does not generalize directly to the Finsler setup, and one needs to look at reversible Finsler metrics, which are symmetric in nature. For submanifolds, geodesics emanating from them and also ending in them, possibly at a point different from the origin, has been studied by Omori in [Omo68]. Generalizing results from [ISS12] and [Xu15], we have the following, which is new even for the Riemannian setup.
Theorem D (4.17).
[…]
The above result follows from 4.15, which provides a criterion for the existence of exactly two distance minimizing geodesics from a manifold to its cut locus and might be of independent interest.
Organization of the paper
In section 2, we begin by recalling basic definitions from Finsler geometry, including the Finsler metric (2.1) and the Chern connection (2.3). These fundamental concepts lay the groundwork for our subsequent discussions. In the next section (section 3), we introduce the cut locus (3.16) and the focal locus (3.22) of a submanifold, and derive some basic properties of them. In section 4, we present our main results including the characterization of the cut locus, and the deformation of its complement. The last part (subsection 4.5) of this section, we address the generalization of Klingenberg’s lemma to the Finsler setup.
2. Preliminaries on Finsler Geometry
This section recalls some basic notions about Finsler geometry and cut locus. Primary references for this section are [She01, Pet06, JS15, Oht21]. Throughout this article, we shall use the notation $\widehat{TM}=TM\setminus 0$ to denote the slit tangent bundle of a manifold $M$. Given a bundle $E$ over a manifold, $\Gamma E$ will be considered as the sheaf of sections of $E$, in particular, $X\in\Gamma E$ denotes a local section of $E$ with some unspecified open set in $M$ as the domain of definition.
2.1. Finsler Metric
Finsler metric was first studied by P. Finsler in his dissertation [Fin51]. In his thesis, he studied the variational problems of Finsler metrics.
Definition 2.1.
Let $M$ be a smooth manifold, and $TM$ denotes its tangent bundle. A Finsler metric on $M$ is a continuous function $F:TM\to\mathbb{R}$ satisfying the following properties.
  1. (1)
    $F$ is smooth on $\widehat{TM}$;
  2. (2)
    for any $p\in TM$, the restriction $F_{p}\vcentcolon=F\big{|}_{T_{p}M}$ is a Minkowski norm, i.e.,
    1. (a)
      (Positive 111-homogeneity) for any $\lambda>0$ and $\mathbf{v}\in T_{p}M\setminus\{0\}$, we have $F_{p}(\lambda\mathbf{v})=\lambda F_{p}(\mathbf{v})$, and
    2. (b)
      (Strong Convexity) for all $\mathbf{v}\in T_{p}M\setminus\{0\}$, the symmetric tensor $g_{\mathbf{v}}$ on $T_{p}M$, called the fundamental tensor, is positive definite, where
      $g_{\mathbf{v}}(\mathbf{v}_{1},\mathbf{v}_{2})\vcentcolon=\left.\dfrac{1}{2}\dfrac{\partial^{2}}{\partial s_{1}\partial s_{2}}\right|_{s_{1}=s_{2}=(0,0)}\left(F_{p}(\mathbf{v}+s_{1}\mathbf{v}_{1}+s_{2}\mathbf{v}_{2})\right)^{2}.$
$F$ is called a pseudo-Finsler (or, semi-Finsler) metric if the fundamental tensor is only nondegenerate. $F$ is reversible if $F(-\mathbf{v})=F(\mathbf{v})$ holds for all $\mathbf{v}\in\widehat{TM}$.
Remark 2.2.
Although we only state them for Finsler manifolds, all the results presented in this article remain true for pseudo-Finsler manifolds, without any significant changes to the proof, unless mentioned otherwise.
As a consequence of the homogeneity, the fundamental tensor satisfies the useful identity
$g_{\mathbf{v}}(\mathbf{v},\mathbf{v})=F_{p}(\mathbf{v})^{2},\quad\mathbf{v}\in T_{p}M\setminus 0.$
For $\mathbf{v}\in T_{p}M\setminus 0$, the associated Cartan tensor on $T_{p}M$ is a symmetric 333-tensor defined as
$C_{\mathbf{v}}(\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3})\vcentcolon=\left.\frac{1}{4}\frac{\partial^{3}}{\partial s_{1}\partial s_{2}\partial s_{3}}\right|_{s_{1}=s_{2}=s_{3}=0}\left(F_{p}(\mathbf{v}+s_{1}\mathbf{v}_{1}+s_{2}\mathbf{v}_{2}+s_{3}\mathbf{v}_{3})\right)^{2}.$
For each $\mathbf{v}\in T_{p}M\setminus 0$ and $\mathbf{u},\mathbf{w}\in T_{p}M$, we have
$C_{\mathbf{v}}(\mathbf{v},\mathbf{u},\mathbf{w})=C_{\mathbf{v}}(\mathbf{u},\mathbf{v},\mathbf{w})=C_{\mathbf{v}}(\mathbf{u},\mathbf{w},\mathbf{v})=0.$
We extend the definition of the fundamental tensor to vector fields as follows:
$g_{V}(X,Y)(p)\vcentcolon=g_{V_{p}}(X_{p},Y_{p}),$
for any $V\in\Gamma\widehat{TM}$ and $X,Y\in\Gamma TM$ defined near $p\in M$. Similarly, we extend the definition of the Cartan tensor as well.
2.1.1. Chern Connection
Unlike the Levi-Civita connection in the Riemannian context, we have a family of torsionless connections on a Finsler manifold.
Definition 2.3.
[Rad04, Jav14] For each $V\in\Gamma\widehat{TM}$, we have a unique affine connection
$\nabla^{V}:\Gamma TM\otimes\Gamma TM\rightarrow\Gamma TM,$
called the Chern connection, satisfying the following conditions for any $X,Y,Z\in\Gamma TM$.
  • (Torsion freeness) $\nabla^{V}_{X}Y-\nabla^{V}_{Y}X=[X,Y]$
  • (Almost metric compatibility) $X(g_{V}(Y,Z))=g_{V}(\nabla^{V}_{X}Y,Z)+g_{V}(Y,\nabla^{V}_{X}Z)+\frac{1}{2}C_{V}(\nabla^{V}_{X}V,Y,Z)$
The associated curvature tensor is given by
$R^{V}(X,Y)Z\vcentcolon=\nabla^{V}_{X}\nabla^{V}_{Y}Z-\nabla^{V}_{Y}\nabla^{V}_{X}Z-\nabla^{V}_{[X,Y]}Z,\quad X,Y,Z\in\Gamma TM,$
which is skewsymmetric: $R^{V}(X,Y)=-R^{V}(Y,X)$.
Definition 2.4.
Given a curve $\gamma:[a,b]\rightarrow M$ and $W\in\Gamma\gamma^{*}\widehat{TM}$, the covariant derivative along $\gamma$ is defined as
$D^{W}_{\gamma}:\Gamma\gamma^{*}TM\rightarrow\Gamma\gamma^{*}TM,$
which satisfies the following metric compatibility
$\frac{d}{dt}g_{W}(X,Y)=g_{W}(D^{W}_{\gamma}X,Y)+g_{W}(X,D^{W}_{\gamma}Y)+2C_{W}(D^{W}_{\gamma}W,X,Y),$
for any $X,Y\in\Gamma\gamma^{*}TM$. If $\dot{\gamma}(t)\neq 0$, then a vector field $X\in\Gamma\gamma^{*}TM$ is said to parallel with respect to $\gamma$ if $D^{\dot{\gamma}}_{\gamma}X=0$.
2.1.2. Legendre Transformation
We have a nonlinear fiber preserving map $\mathfrak{L}:\widehat{TM}\rightarrow\widehat{T^{*}M}$ given by
$\mathfrak{L}(\mathbf{v})(\mathbf{w})\vcentcolon=g_{\mathbf{v}}(\mathbf{v},\mathbf{w}),\quad\mathbf{v}\in T_{p}M\setminus 0,\;\mathbf{w}\in T_{p}M.$
[…]
2.2. Geodesics and Jacobi Fields
Let us denote the space of piecewise smooth paths $\gamma:[a,b]\rightarrow M$ as $\mathcal{P}=\mathcal{P}([a,b])$. We have two functionals defined on this path space.
  • (Length Functional)
    $L:\mathcal{P}$$\rightarrow\mathbb{R}$
[…]
  • (Energy Functional)
    $E:\mathcal{P}$$\rightarrow\mathbb{R}$
[…]
Note that $L$ is independent of the reparametrization of a curve, whereas $E$ is not. A geodesic is defined as a critical point of the energy functional in the variational sense.
Let us recall the notion of variation of a given curve. Given a $\gamma\in\mathcal{P}$, we identify the tangent space to $\mathcal{P}$ at $\gamma$ as the infinite dimensional vector space of vector fields along $\gamma$:
$T_{\gamma}\mathcal{P}\coloneqq\Gamma\gamma^{*}TM.$
Given a vector field $W\in\Gamma\gamma^{*}TM$, a $W$-variation of $\gamma$ is a piecewise smooth map $\Lambda:(-\epsilon,\epsilon)\times[a,b]\rightarrow M$ satisfying
$\Lambda(0,t)=\gamma(t),\quad\left.\frac{\partial}{\partial s}\right|_{s=0}\Lambda(s,t)=W(t),$
for all $(s,t)$ in the domain. We tacitly assume that each curve $t\mapsto\Lambda_{s}(t)=\Lambda(s,t)$ has the same breakpoints as $\gamma$. Lemma 2.1 from Sakai’s book [Sak96] can be easily generalized to the Finsler setup to prove the existence of $W$-variation.
Definition 2.5.
A variation $\Lambda$ of a curve $\gamma$ joining $p=\gamma(a)$ to $q=\gamma(b)$ is said to be proper if the endpoints are kept fixed, i.e.,
$\Lambda(s,a)=p,\quad\Lambda(s,b)=q,\qquad-\epsilon<s<\epsilon.$
Similarly, a vector field $W\in\Gamma\gamma^{*}TM$ is said to be proper if $W(a)=0=W(b)$ holds.
We refer to [Oht21, Chapter 7] for the first and second variation formulas for the Length functional (with respect to proper variations), and as for the energy functional we refer to [JS15].
Definition 2.6.
A piecewise $C^{1}$ curve $\gamma:[a,b]\rightarrow M$ is called a geodesic if it is a critical point of the energy functional with respect to proper variations, i.e., for any proper variation $\Lambda$ of $\gamma$ we have $\left.\frac{d}{ds}\right|_{s=0}E(\Lambda_{s})=0$.
As usual, a geodesic is locally distance minimizing, where we have the following notion of Finsler distance.
Definition 2.7.
Given a Finsler manifold $(M,F)$, the distance between points $p,q\in M$ is defined as
$d(p,q)=\inf\left\{L(\gamma)\;\middle|\;\gamma\in\mathcal{P},\gamma(a)=p,\gamma(b)=q\right\}.$
Remark 2.8.
In general, the distance function is asymmetric, although it induces the same topology as that of the manifold. If $F$ is reversible, then $d$ is a symmetric metric (see, e.g., [BCS00, pg. 151]).
Proposition 2.9.
Given a unit-speed curve $\gamma$ (i.e., $F(\dot{\gamma}(t))=1$ for all $t$), the following are equivalent.
  • $\gamma$ is a geodesic.
  • $\gamma$ is a critical point of the length functional with respect to proper variations.
  • $\gamma$ is locally distance minimizing, i.e., for any $a<t_{0}<b$, there exists an open interval containing $t_{0}$, say, $t_{0}\in(a_{t_{0}},b_{t_{0}})\subset[a,b]$, such that $d(\gamma(s),\gamma(t))=t-s$ holds for any $a_{t_{0}}\leq s\leq t\leq b_{t_{0}}$.
  • $\dot{\gamma}$ is a parallel vector field with respect to $\gamma$, i.e., $D^{\dot{\gamma}}_{\gamma}\dot{\gamma}=0$.
For proof of the above equivalences, we refer to [Oht21]. Geodesics are precisely the solutions to an initial value problem known as the geodesic equation (e.g., [Oht21, Equation 3.15]). As such, geodesics are always smooth, and given a vector $\mathbf{v}\in T_{p}M$, we have a unique maximal geodesic $\gamma_{\mathbf{v}}:[0,\ell]\rightarrow M$ satisfying
$\gamma(0)=p,\qquad\dot{\gamma}(0)=\mathbf{v}.$
Note that, unlike Riemannian geometry, due to the asymmetry of the Finsler metric, a geodesic traversed in the reverse direction need not be a geodesic. Even then, the following observation remains true.
Proposition 2.10.
Two distinct geodesics in a Finsler manifold always intersect transversally.
Proof.
Let $\gamma:[0,a]\rightarrow M$ and $\eta:[0,b]\rightarrow M$ be two distinct geodesics meeting at $p=\gamma(a)=\eta(b)$. If possible, suppose $\mathbf{v}=\dot{\gamma}(a)=\dot{\eta}(b)$. Then, the reversed curves $\bar{\gamma}:[a,0]\rightarrow M,\bar{\eta}:[b,0]\rightarrow M$ are geodesics for the reverse Finsler metric ([Oht21, Section 2.5]), with same initial velocity $-\mathbf{v}$. Consequently, they are identical. But then $\gamma,\eta$ are identical as well, which is a contradiction. ∎
By the Hopf-Rinow theorem ([Oht21, Theorem 3.21]), a Finsler manifold $(M,F)$ is forward (resp. backward) complete if and only if given any two points $p,q\in M$ there exists a geodesic joining $p$ to $q$ (resp. $q$ to $p$). In particular, for a forward (or backward) complete Finsler manifold, the domain of definition of $\gamma_{\mathbf{v}}$ can be taken to be [0,)[0$\infty$)[ 0 , ∞ ). Throughout this article, we shall assume $(M,F)$ to be a forward complete Finsler manifold, unless otherwise mentioned.
Definition 2.11.
The exponential map
$\exp:TM\rightarrow M$
at a point $p\in M$ is defined as $\exp_{p}(\mathbf{v})=\gamma_{\mathbf{v}}(1)$ for any $\mathbf{v}\in T_{p}M$, where $\gamma_{\mathbf{v}}:[0,\infty)\rightarrow M$ is the unique geodesic starting at $p$ in the direction of $\mathbf{v}$.
It follows from the theory of ordinary differential equation that the exponential map is smooth on $\widehat{TM}$, but only $C^{1}$ on the whole tangent bundle. In fact, the exponential is $C^{2}$ on $TM$ precisely when the Finsler metric comes from a Riemannian metric, in which case the exponential map is $C^{\infty}$ on $TM$ as well.
Definition 2.12.
A geodesic $\gamma:[0,\ell]\rightarrow M$ is said to be a global distance minimizer (or simply a minimizer) if $d(\gamma(0),\gamma(\ell))=\ell=L(\gamma)$.
We have the following easy observation.
Proposition 2.13.
If $\gamma:[0,\ell]\rightarrow M$ is a minimizer, then for any $0\leq a\leq b\leq\ell$ we have $\gamma|_{[a,b]}$ is a minimizer, and in particular $d(\gamma(a),\gamma(b))=b-a$.
Any piecewise smooth curve $\gamma$ joining $p$ to $q$, and satisfying $L(\gamma)=d(p,q)$ is, up to a reparametrization, a minimizing geodesic from $p$ to $q$, and in particular $\gamma$ is smooth. A consequence of the completeness assumption is that any two points $p,q\in M$ can be joined by a minimizer, say, $\gamma:[0,\ell]\rightarrow M$ so that
$\gamma(0)=p,\qquad\gamma(\ell)=q,\qquad d(p,q)=\ell=L(\gamma).$
2.2.1. Jacobi Fields
Definition 2.14.
Given a unit-speed geodesic $\gamma:[a,b]\rightarrow M$, a variation $\Lambda:(-\epsilon,\epsilon)\times[a,b]\rightarrow M$ of $\gamma$ is said to be a geodesic variation, if for each $-\epsilon<s<\epsilon$, the curve $\Lambda_{s}:[a,b]\rightarrow M$ is a constant-speed geodesic.
Definition 2.15.
Let $\gamma:[a,b]\rightarrow M$ be a unit-speed geodesic. A vector field $J\in\Gamma\gamma^{*}TM$ along $\gamma$ is called a Jacobi field of $\gamma$ if there exists a geodesic variation $\Lambda:(-\epsilon,\epsilon)\times[a,b]\rightarrow M$ of $\gamma$ satisfying
$\left.\frac{\partial}{\partial s}\right|_{s=0}\Lambda(s,t)=J(t),\quad t\in[a,b].$
Proposition 2.16.
Given a unit-speed geodesic $\gamma:[a,b]\rightarrow M$, a vector field $J\in\Gamma\gamma^{*}TM$ is a Jacobi field if and only if it satisfies the Jacobi equation
$D^{\dot{\gamma}}_{\gamma}D^{\dot{\gamma}}_{\gamma}J-R^{\dot{\gamma}}(\dot{\gamma},J)\dot{\gamma}=0.$
Since the Jacobi equation is a second order ODE, given the initial data, $\mathbf{u},\mathbf{v}\in T_{\gamma(a)}M$, there exists a unique Jacobi field $J$ along $\gamma$ satisfying, $J(a)=\mathbf{u}$ and $D^{\dot{\gamma}}_{\gamma}J(a)=\mathbf{v}$. In particular, the collection of all Jacobi fields along $\gamma$ forms a vector space of dimension $2\dim M$.
Proposition 2.17.
Let $\mathbf{v}\in T_{p}M$ with $F(\mathbf{v})=1$, and $\gamma(t)=\exp_{p}(t\mathbf{v})$ be the geodesic in the direction of $\mathbf{v}$. Then, the kernel of the map
$d(\exp_{p})|_{t\mathbf{v}}:T_{t\mathbf{v}}(T_{p}M)=T_{p}M\rightarrow T_{\gamma(t)}M$
is naturally isomorphic to the vector space
$\left\{J\;\middle|\;\text{$J$ is a Jacobi field along $\gamma$, satisfying $J(0)=0=J(t)$.}\right\}.$
3. Submanifolds in Finsler Manifold
In this section, we generalize the notions from the previous section.
3.1. Normal Cone Bundles
Given a submanifold $N$, it is natural to ask for a notion of a normal bundle to it.
Definition 3.1.
Given a submanifold $N\subset M$, the set
$\nu_{p}=\nu_{p}(N)=\left\{\mathbf{v}\in T_{p}M\;|\;g_{\mathbf{v}}(\mathbf{v},\mathbf{w})=0\;\forall\mathbf{w}\in T_{p}N\right\}$
is called the normal cone of $N$ at $p\in N$. The set $\nu=\nu(N)=\cup_{p\in N}\nu_{p}(N)$ is called the normal cone bundle of $N$.
It should be noted that $\nu(N)$ is not a vector bundle in general; in fact, it is a cone bundle. Nevertheless, this is the obvious replacement for the notion of a normal bundle as in the Riemannian setting. We shall denote $\hat{\nu}_{p}\vcentcolon=\nu_{p}\setminus\{0\}$, and $\hat{\nu}\vcentcolon=\cup_{p\in N}\hat{\nu}_{p}$ is then the slit cone bundle. The Legendre transformation $\mathfrak{L}$ maps $\nu$ homeomorphically onto the annihilator bundle $\nu^{*}\subset T^{*}M|_{N}$ of $TN$ defined as
$\nu^{*}|_{p}\vcentcolon=\left\{\omega\in T_{p}^{*}M\;\middle|\;\omega|_{T_{p}N}=0\right\},\quad p\in N.$
It follows that $\hat{\nu}$ is a smooth submanifold of $TM$ of dimension $\dim N+\operatorname{codim}N=\dim M$, whereas $\nu$ is only a topological submanifold. Note that $\nu$ is a closed subset of $TM$.
Definition 3.2.
The unit cone bundle of $N$ is denoted as $S(\nu)=\cup_{p\in N}S(\nu_{p})$, where
$S(\nu_{p})\vcentcolon=\left\{\mathbf{v}\in\nu_{p}\;\middle|\;F_{p}(\mathbf{v})=1\right\}.$
Remark 3.3.
If $N$ is a compact submanifold, then $S(\nu)$ is again a compact submanifold of $TM$ of dimension $\dim M-1$.
Definition 3.4.
Given $N\subset M$, the normal exponential map
$\exp^{\nu}:\nu(N)\rightarrow M$
is defined as the restriction of the exponential map to the cone bundle $\nu(N)$.
As $\hat{\nu}$ is a smooth submanifold of $TM$, the restricted map $\exp^{\nu}|_{\hat{\nu}}$ is smooth.
3.2. $N$-Geodesics and $N$-Jacobi Fields
Given a submanifold $N\subset M$, we have the subspace of piecewise smooth paths
$\mathcal{P}_{N}=\mathcal{P}_{N}([a,b])=\{\gamma\in\mathcal{P}([a,b])\;|\;\gamma(a)\in N\}$
starting at $N$. For any $\gamma\in\mathcal{P}_{N}$ we say $\gamma$ is a curve joining $N$ to $\gamma(b)$.
Definition 3.5.
Given $q\in M$, the distance from $N$ to $q$ is defined as
$d(N,q)\vcentcolon=\inf\left\{L(\gamma)\;\middle|\;\gamma\in\mathcal{P}_{N},\;\gamma(b)=q\right\}.$
We identify the tangent space of $\mathcal{P}_{N}$ at a given $\gamma\in\mathcal{P}_{N}$ as the infinite dimensional vector space
$T_{\gamma}\mathcal{P}_{N}\vcentcolon=\left\{W\in\Gamma\gamma^{*}TM\;\middle|\;W(a)\in T_{\gamma(a)}N\right\}.$
Given $W\in T_{\gamma}\mathcal{P}_{N}$, a $W$-variation is a piecewise smooth map $\Lambda:(-\epsilon,\epsilon)\times[a,b]\rightarrow M$ satisfying
$\Lambda(0,t)=\gamma(t),\quad\left.\frac{\partial}{\partial s}\right|_{s=0}\Lambda(s,t)=W(t),\quad\Lambda(s,a)\in N,$
for all $(s,t)$ in the domain of the definition. The variation is said to be proper if it further satisfies $\Lambda(s,b)=\gamma(b)$ for all $-\epsilon<s<\epsilon$. Clearly, proper variations correspond to those $W\in T_{\gamma}\mathcal{P}_{N}$ that satisfy $W(b)=0$. This leads to the first and second variational formulas for the restricted length and energy functionals: $L|_{\mathcal{P}_{N}}$ and $E|_{\mathcal{P}_{N}}$. We refer to [JS15] for the relevant formulas.
Definition 3.6.
A piecewise $C^{1}$ curve $\gamma:[a,b]\rightarrow M$ in $\mathcal{P}(N)$ is called an $N$-geodesic if $\gamma$ is a critical point of the restricted energy functional $E|_{\mathcal{P}(N)}$ with respect to proper variations, i.e., for any proper variation $\Lambda$ of $\gamma$ we have $\left.\frac{d}{ds}\right|_{s=0}E(\Lambda_{s})=0$
Proposition 3.7.
Given a unit-speed curve $\gamma:[a,b]\rightarrow M$ in $\mathcal{P}(N)$, the following are equivalent.
  • $\gamma$ is an $N$-geodesic
  • $\gamma$ is a critical point of the restricted length functional $L|_{\mathcal{P}(N)}$ with respect to proper variations.
  • $\gamma(t)=\exp^{\nu}(t\mathbf{v})$, where $p=\gamma(0)$ and $\mathbf{v}=\dot{\gamma}(a)\in S(\nu_{p})$.
For proof of the above equivalences, we refer to [Zha17]. Next, we define the notion of global distance minimizer from $N$.
Definition 3.8.
An $N$-geodesic $\gamma:[0,\ell]\to M$ is called an $N$-segment joining $N$ to $\gamma(\ell)$ if $d(N,\gamma(\ell))=L(\gamma)=\ell$.
Proposition 3.9.
Suppose $(M,F)$ is a forward complete Finsler manifold, and $N$ is a closed submanifold of $M$. Then, for any $q\in M$ there exists an $N$-segment joining $N$ to $q$.
Similar to 2.13, we have the following result.
Proposition 3.10.
An $N$-geodesic $\gamma:[0,\ell]\rightarrow M$ is an $N$-segment if and only if $d(N,\gamma(t))=t$ for all $0\leq t\leq\ell$.
Proof.
Suppose $\gamma$ is an $N$-segment. Then, $\gamma$ is in particular a minimizer joining $\gamma(0)$ to $\gamma(\ell)$. If possible, say, there exists some $0<T<\ell$ so that $d(N,\gamma(T))<T=L(\gamma|_{[0,T]})$. Then, there exists some $N$-segment $\eta:[0,T_{0}]\rightarrow M$ joining $N$ to $q=\gamma(T)=\eta(T_{0})$, so that $L(\eta)=d(N,\gamma(T))<L(\gamma|_{[0,T]})$. Consider the broken curve $\zeta$ obtained by concatenating $\eta$ and $\gamma|_{[T,\ell]}$. It follows that
$L(\zeta)=L(\eta)+L(\gamma|_{[T,\ell]})<L(\gamma|_{[0,T]})+L(\gamma|_{[T,\ell]})=L(\gamma)=d(N,\gamma(\ell)),$
which contradicts the fact that $\zeta$ is a curve joining $N$ to $\gamma(\ell)$. The converse statement is obvious. ∎
3.2.1. $N$-Jacobi Fields
Definition 3.11.
Given a unit-speed $N$-geodesic $\gamma:[a,b]\rightarrow M$, a variation $\Lambda:(-\epsilon,\epsilon)\times[a,b]\rightarrow M$ is said to be an $N$-geodesic variation if for each $-\epsilon<s<\epsilon$, the curve $\Lambda_{s}:[a,b]\rightarrow M$ is a constant speed $N$-geodesic.
Definition 3.12.
Let $\gamma:[a,b]\rightarrow M$ be a unit-speed $N$-geodesic. A vector field $J\in\Gamma\gamma^{*}TM$ is called an $N$-Jacobi field of $\gamma$ if there exists an $N$-geodesic variation $\Lambda:(-\epsilon,\epsilon)\times[a,b]\rightarrow M$ satisfying
$\left.\frac{\partial}{\partial s}\right|_{s=0}\Lambda(s,t)=J(t),\quad t\in[a,b].$
Proposition 3.13.
[JS15] Given a unit-speed $N$-geodesic $\gamma:[a,b]\rightarrow M$, a vector field $J\in\Gamma\gamma^{*}TM$ is an $N$-Jacobi field if and only if it satisfies
$D^{\dot{\gamma}}_{\gamma}D^{\dot{\gamma}}_{\gamma}J-R^{\dot{\gamma}}(\dot{\gamma},J)\dot{\gamma}=0,\quad J(a)\in T_{\gamma(a)}N,\quad\pi D^{\dot{\gamma}}_{\gamma}J(a)=\pi\nabla^{\dot{\gamma}}_{J}\dot{\gamma}(a),$
where $\pi$ is the first component projection of the canonical splitting $T_{\gamma(a)}M=T_{\gamma(a)}N\oplus\left(T_{\gamma(a)}N\right)^{\perp_{g_{\dot{\gamma}(a)}}}$.
From the second variation formula of the restricted energy functional $E|_{\mathcal{P}_{N}}$, we get the index form.
Definition 3.14.
Let $\gamma:[a,b]\rightarrow M$ be a unit-speed $N$-geodesic. Then, the index form $I_{\gamma}$ is defined for $X,Y\in T_{\gamma}\mathcal{P}_{N}$ with $X(b)=0=Y(b)$ as
$\mathcal{I}_{\gamma}(X,Y)\coloneqq\int_{a}^{b}\left[g_{\dot{\gamma}}(D^{\dot{\gamma}}_{\gamma}X,D^{\dot{\gamma}}_{\gamma}Y)-g_{\dot{\gamma}}\left(R^{\dot{\gamma}}(\dot{\gamma},X)Y,\dot{\gamma}\right)\right]-g_{\dot{\gamma}(a)}\left(\pi^{\perp}\nabla^{\dot{\gamma}(a)}_{X}Y,\dot{\gamma}(a)\right),$
where $\pi^{\perp}$ is the second component projection of the canonical splitting $T_{\gamma(a)}M=T_{\gamma(a)}N\oplus\left(T_{\gamma(a)}N\right)^{\perp_{g_{\dot{\gamma}(a)}}}$.
It follows that $I_{\gamma}$ is a symmetric 222-form, and the kernel of the index form $I_{\gamma}$ consists of precisely the $N$-Jacobi fields along $\gamma$.
3.3. Cut Locus of a Submanifold
Definition 3.15.
The cut locus of a point $p\in M$ is the set consisting of all $q\in M$ such that there exists a minimizer from $p$ to $q$, any extension of which fails to be distance minimizing. We denote the cut locus of $p$ by $\mathrm{Cu}(p)$.
Generalizing the notion of a cut locus of a point to an arbitrary submanifold (or, more generally, any subset), we have the following definition.
Definition 3.16.
Given a submanifold $N\subset M$, the cut locus of $N$ consists of points $q\in M$ such that there exists an $N$-segment joining $N$ to $q$, whose extension fails to be an $N$-segment.
Proposition 3.17.
$q\in\mathrm{Cu}(N)$ if and only if for any $N$-segment $\gamma$ joining $N$ to $q$, an extension of $\gamma$ fails to be an $N$-segment.
Proof.
Suppose $q\in\mathrm{Cu}(N)$. We then have an $N$-segment $\gamma:[0,\ell]\rightarrow M$ with $\ell=d(N,q)$ and $q=\gamma(\ell)$, whose extension fails to be an $N$-segment. If possible, let $\eta:[0,\ell]\rightarrow M$ be another $N$-segment joining $N$ to $q=\eta(\ell)$, which remains an $N$-segment when extended to $\eta:[0,\ell+\epsilon]\rightarrow M$ for some $\epsilon>0$. Since $\gamma,\eta$ are distinct, we have $\dot{\gamma}(\ell)\neq\dot{\eta}(\ell)$ (2.10). Thus, we have a broken curve $\zeta$ obtained from concatenating $\gamma$ and $\eta|_{[\ell,\ell+\epsilon]}$. Now, $L(\zeta)=L(\gamma)+L(\eta|_{[\ell,\ell+\epsilon]})=\ell+\epsilon=d(N,\eta(\ell+\epsilon))$, i.e., $\zeta$ achieves the minimal distance. But then $\zeta$ must be a smooth curve, which is a contradiction.
The converse statement is obvious. ∎
Definition 3.18.
Given $\mathbf{v}\in S(\nu)$, the cut time of $\mathbf{v}$ is defined as
$\rho(\mathbf{v})\vcentcolon=\sup\left\{t\;\middle|\;d(N,\gamma_{\mathbf{v}}(t\mathbf{v}))=t\right\},$
where $\gamma_{\mathbf{v}}(t)=\exp^{\nu}(t\mathbf{v})$ is the unique $N$-geodesic with initial velocity $\mathbf{v}$.
Note that we allow $\rho$ to take the value $\infty$, although if $M$ is compact then $\rho(\mathbf{v})<\infty$ for any $\mathbf{v}\in\nu$. Indeed, we shall see in 4.7 that the map $\rho:\nu\rightarrow[0,\infty]$ is continuous, and consequently, $\rho$ is finite if $N$ is closed. It follows from [AJ19] that $\rho$ is always strictly positive.
Definition 3.19.
Given a submanifold $N\subset M$, the tangent cut locus of $N$ is defined as
$\widetilde{\mathrm{Cu}}(N)\vcentcolon=\left\{\rho(\mathbf{v})\mathbf{v}\;\middle|\;\mathbf{v}\in S(\nu),\;\rho(\mathbf{v})\neq\infty\right\}\subset\nu.$
It follows from 3.16 that $\mathrm{Cu}(N)=\exp^{\nu}(\widetilde{\mathrm{Cu}}(N))$.
3.4. Focal Locus of a Submanifold
Definition 3.20.
Given a submanifold $N\subset M$, a vector $\mathbf{v}\in\hat{\nu}$ is said to be a tangent focal point of $N$ if
$d(\exp^{\nu})|_{\mathbf{v}}:T_{\mathbf{v}}\hat{\nu}\rightarrow T_{\exp^{\nu}(\mathbf{v})}M$
is degenerate, i.e., if $\mathbf{v}$ is a critical point of $\exp^{\nu}|_{\hat{\nu}}$. The nullity of the map is known as the index of the tangent focal point $\mathbf{v}$. The collection of all such vectors is called the tangent focal locus of $N$.
Definition 3.21.
Given $N\subset M$ and $\mathbf{v}\in S(\nu)$, we define the (first) focal time as
$\lambda(\mathbf{v})=\inf\left\{t\;\middle|\;\text{$d(\exp^{\nu})|_{t\mathbf{v}}$ is degenerate}\right\}.$
We set $\lambda(\mathbf{v})=\infty$ if there are no focal points of $N$ in the direction of $\mathbf{v}$.
Definition 3.22.
The focal locus of $N$ is defined as the image of the tangent focal locus under the $\exp^{\nu}$ map. In other words, the focal locus of $N$ consists of the critical values of the (restricted) normal exponential map $\exp^{\nu}|_{\hat{\nu}}$. The first focal locus of $N$ is defined as the set $\left\{\exp^{\nu}(\lambda(\mathbf{v})\mathbf{v})\mid\mathbf{v}\in S(\nu),\lambda(\mathbf{v})\neq\infty\right\}$.
In the literature, the focal locus of a submanifold (of a Riemannian manifold) is also known as the conjugate locus. In this article, we reserve the term conjugate locus for when $N$ is a point.
Proposition 3.23.
Let $\mathbf{v}\in S(\nu)$ and $\gamma$ be the unit-speed $N$-geodesic given as $\gamma(t)=\exp^{\nu}(t\mathbf{v})$. Then for any $t>0$, the kernel of the map
$d\left(\exp^{\nu}|_{\hat{\nu}}\right)\big{|}_{t\mathbf{v}}:T_{t\mathbf{v}}\hat{\nu}\rightarrow T_{\gamma(t)}M$
is naturally isomorphic to the vector space
$\left\{J\;\middle|\;\text{$J$ is an $N$-Jacobi field along $\gamma$, satisfying $J(t)=0$}\right\}.$
4. Main Results
We start from an important characterization of the cut locus in terms of focal locus. A result of Basu and Prasad [BP21] may be generalized to prove that the separating set is dense in the cut locus (4.8). We will also see that the complement of the cut locus and submanifold strongly deforms to the submanifold and the cut locus, respectively. We will prove the existence of $N$-geodesic loop which is a generalization of the result of [ISS12].
4.1. Characterization of Cut Locus
In this section, we discuss two ways one can characterize the cut locus of a submanifold, which are well-known in the Riemannian setting. Recall first the definition.
Definition 4.1.
Given a submanifold $N\subset M$, a point $p\in M$ is said to be a separating point of $N$ if there exists two distinct $N$-segments joining $N$ to $p$. The collection of all separating points of $N$ is called the separating set (or the Maxwell strata) of $N$, denoted $\mathrm{Se}(N)$,
Proposition 4.2.
[…]
Proof.
Let $q\in\mathrm{Se}(N)$, and take two $N$-segments $\gamma_{1},\gamma_{2}:[0,\ell]\to M$ joining, respectively, $p_{1},p_{2}\in N$ to $q$. If possible, suppose $q\not\in\mathrm{Cu}(N)$. Then, a sufficiently small extension of $\gamma_{2}$, say, $\gamma_{2}:[0,\ell+\epsilon]\to M$ remains an $N$-segment (3.17), and hence, $d(N,\gamma_{2}(\ell+\epsilon))=\ell+\epsilon$.
$\gamma_{1}$
$\gamma_{2}$
$N$
$p_{1}$
$p_{2}$
$\gamma_{1}(\ell)=q=\gamma_{2}(\ell)$
$\gamma_{2}(\ell+\epsilon)$
$\eta$
Figure 1. Constructing the broken curve $\eta$.
On the other hand, by 2.10, $\dot{\gamma}_{1}(\ell)\neq\dot{\gamma}_{2}(\ell)$, and so we have a broken curve $\eta:[0,\ell+\epsilon]\to M$ obtained from concatenating $\gamma_{1}$ and $\gamma_{2}|_{[\ell,\ell+\epsilon]}$ (Figure 1). In particular, from the triangle inequality, we get
$d(p_{1},\eta(\ell+\epsilon))<d(p_{1},\eta(\ell))+d(\eta(\ell),\eta(\ell+\epsilon))=\ell+\epsilon.$
Consequently,
$d(N,\gamma_{2}(\ell+\epsilon))=d(N,\eta(\ell+\epsilon))\leq d(p_{1},\eta(\ell+\epsilon))<\ell+\epsilon.$
This is a contradiction and so, $q\in\mathrm{Cu}(N)$. ∎
In fact, for any point on an $N$-segment not in the cut locus, we have a uniqueness.
Proposition 4.3.
[…]
Proof.
Suppose, if possible, there exists some $N$-segment $\eta:[0,t_{0}]\to M$ joining $N$ to $q_{0}=\gamma(t_{0})$, which is distinct from $\gamma|_{[0,t_{0}]}$. Then $q_{0}\in\mathrm{Se}(N)\subset\mathrm{Cu}(N)$ by 4.2. On the other hand, $\gamma$ remains an $N$-segment beyond $\gamma(t_{0})$, and so $q_{0}\not\in\mathrm{Cu}(N)$ by 3.17. This is a contradiction. ∎
Let us now recall the following useful result, which is known for the case when $N$ is a point ([Oht21, Prop 7.16, pg. 89]).
Lemma 4.4.
Let $\gamma:[0,\ell]\to M$ be a unit speed $N$-geodesic. If $\gamma(T)$ is a focal point of $N$ along $\gamma$ for some $T\in(0,\ell)$, then $d(N,\gamma(t))<t$ for all $t\in(T,\ell]$. In particular, if $\gamma|_{[0,T]}$ is an $N$-segment, then $\gamma(T)\in\mathrm{Cu}(N)$.
Proof.
Without loss of generality, we may assume that $\gamma(T)$ is the first focal point of $N$ along $\gamma$. Let $J_{1}$ be a non-zero $N$-Jacobi field along $\gamma$ such that $J_{1}(T)=0$ and $J_{1}(0)\in T_{p}N$, where $p=\gamma(0)$. We extend $J_{1}$ to $[0,\ell]$ by 0. By a similar argument as in [Sak96, Lemma 4.9, pg. 61], we can find an $\epsilon>0$ so that $\gamma|_{[T-\epsilon,T+\epsilon]}$ has no pair of conjugate points. Let us consider the (unique) Jacobi field $J_{2}$ along $\gamma|_{[T-\epsilon,T+\epsilon]}$ satisfying $J_{2}(T-\epsilon)=J_{1}(T-\epsilon)$ and $J_{2}(T+\epsilon)=0$. Since $\gamma|_{[T-\epsilon,T+\epsilon]}$ has no pair of conjugate points, $J_{2}\neq 0$ on $[T-\epsilon,T+\epsilon)$. Next, consider the piecewise smooth vector field $J$ along $\gamma$ defined as follows
$J(t)=\begin{cases}J_{1}(t),&t\in[0,T-\epsilon]\\ J_{2}(t),&t\in[T-\epsilon,T+\epsilon]\\ 0,&t\in[T+\epsilon,\ell].\end{cases}$
[…]
Now, suppose $\gamma_{s}:[0,T]\rightarrow M$ is a $J$-variation of $\gamma$, so that $\gamma_{0}=\gamma$. Since $\gamma$ is a geodesic, we have $E^{\prime}(0)=\frac{d}{ds}|_{s=0}E(\gamma_{s})=0$, and then it follows that $E^{\prime\prime}(0)=\frac{d^{2}}{ds^{2}}|_{s=0}E(\gamma_{s})=I(J,J)<I(J_{1},J_{1})=0$. Hence, for any $s\neq 0$ sufficiently near 0, we have $E(\gamma_{s})<E(\gamma_{0})$. But then $\gamma$ cannot be an $N$-segment, as it is not a local minima of the energy functional. In particular, $\gamma$ fails to be distance minimizing beyond $T$. ∎
Recall that we have $\rho$ as the cut time (3.18), and $\lambda$ as the focal time (3.21) map. Then, as an immediate corollary, we have the following.
Corollary 4.5.
$\rho(\mathbf{v})\leq\lambda(\mathbf{v})$ for all $\mathbf{v}\in S(\nu)$.
The next result is well-known in the Riemannian setting ([BP21]).
Theorem 4.6.
[…]
  1. (1)
    There exists two distinct $N$-segments joining $N$ to $q$, i.e, $q\in\mathrm{Se}(N)$, or
  2. (2)
    $q$ is the first focal point of $N$ along $\gamma$.
Proof.
Suppose $q\in\mathrm{Se}(N)$. Then by 4.2, we have $q\in\mathrm{Cu}(N)$. On the other hand, suppose $q$ is the first focal point of $N$ along $\gamma$, and then by 4.4, any extension of $\gamma$ fails to be distance minimizing from $N$. As $\gamma$ is an $N$-segment, we have $q\in\mathrm{Cu}(N)$.
In order to prove the converse, let $q\in\mathrm{Cu}(N)\setminus\mathrm{Se}(N)$. Then, there exists a unique $N$-segment $\gamma:[0,\ell]\to M$ joining $p=\gamma(0)\in N$ to $q=\gamma(\ell)$, where $\ell=d(N,q)$, so that any extension of $\gamma$ fails to be distance minimizing from $N$. Say, $\mathbf{v}=\dot{\gamma}(0)$, and then $\gamma(t)=\exp^{\nu}(t\mathbf{v})$. Let $q_{i}=\gamma\left(\ell+\frac{1}{i}\right)$. Then, $q_{i}\to q$ and $d(N,q_{i})<L\left(\gamma|_{\left[0,\ell+\frac{1}{i}\right]}\right)$. Let $\gamma_{i}:[0,\ell_{i}]\to M$ be $N$-segments joining $p_{i}=\gamma_{i}(0)\in N$ to $q_{i}=\gamma_{i}(\ell_{i})$, where $\ell_{i}=d(N,q_{i})$ (see Figure 2). Note that $F_{p_{i}}(\mathbf{v}_{i})=1$ where $\mathbf{v}_{i}=\dot{\gamma}_{i}(0)\in\nu_{p_{i}}$, i.e., $(p_{i},v_{i})\in S(\nu)$, the unit normal cone bundle. As $\gamma|_{\left[0,\ell+\frac{1}{i}\right]}$ is not an $N$-segment, $\gamma_{i}$ is distinct from $\gamma$ and in particular, $\mathbf{v}_{i}\neq\mathbf{v}$. Since $N$ is compact, $S(\nu)$ is compact by 3.3. Passing to a subsequence if necessary, we may assume $(p_{i},\mathbf{v}_{i})\rightarrow(p^{\prime},\mathbf{v}^{\prime})\in S(\nu)$.
$p$
$\gamma$
$\gamma(\ell)=q$
$N$
$\mathbf{v}$
$\gamma(\ell+\frac{1}{i})=q_{i}$
$p_{i}$
$\mathbf{v}_{i}$
$\gamma_{i}$
$\gamma\left(\ell_{i}\right)$
Figure 2. Characterization of $\mathrm{Cu}(N)$
Now, consider the $N$-geodesic $\eta(t)=\exp^{\nu}(t\mathbf{v}^{\prime})$. We have
$\ell=d(N,q)=\lim d(N,q_{i})=\lim d(p_{i},q_{i})=\lim\ell_{i}.$
Then, it follows from the continuity of the exponential map that,
$q=\lim q_{i}=\lim\gamma_{i}(\ell_{i})=\lim\exp^{\nu}(\ell_{i}\mathbf{v}_{i})=\exp^{\nu}(\ell\mathbf{v}^{\prime})=\eta(\ell).$
Thus, $\eta$ is an $N$-segment joining $p^{\prime}$ to $q$. Since $q\not\in\mathrm{Se}(N)$, we must have $\gamma=\eta$, and in particular, $(p_{i},\mathbf{v}_{i})\to(p^{\prime},\mathbf{v}^{\prime})=(p,\mathbf{v})$. On the other hand
$\exp^{\nu}(\ell_{i}\mathbf{v}_{i})=\gamma_{i}(\ell_{i})=q_{i}=\exp^{\nu}\left(\left(\ell+\frac{1}{i}\right)\mathbf{v}\right).$
As $\mathbf{v}_{i}\neq\mathbf{v}$, we have $\ell_{i}\mathbf{v}\neq\left(\ell+\frac{1}{i}\right)\mathbf{v}$. Then, the normal exponential map fails to be injective in any neighborhood of $\ell\mathbf{v}$, since both sequences $\ell_{i}\mathbf{v}_{i}\rightarrow\mathbf{v}$ and $\left(\ell+\frac{1}{i}\right)\mathbf{v}\rightarrow\mathbf{v}$. As a consequence of the inverse function theorem, we then have $d(\exp^{\nu})|_{\ell\mathbf{v}}$ is degenerate, i.e., $q=\exp^{\nu}(\ell\mathbf{v})$ is a focal point of $N$ along $\gamma$.
As $\gamma$ is an $N$-segment joining $N$ to $q$, it follows from 4.4 that $\gamma(t)$ is not a focal point of $N$ along $\gamma$ for $0\leq t<T$. Thus, $q$ is the first focal point of $N$ along $\gamma$. This completes the proof. ∎
4.2. Continuity of the $\rho$ and $\lambda$ maps
Recall the cut time map $\rho$ from 3.18. We have the following result.
Theorem 4.7.
The map $\rho:S(\nu)\to[0,\infty]$ is continuous.
Proof.
Consider a converging sequence $(p_{i},\mathbf{v}_{i})\to(p,\mathbf{v})$ in $S(\nu)$. We show that $\rho(p_{i},\mathbf{v}_{i})\to\rho(p,\mathbf{v})$. Let $\ell_{i}=\rho(p_{i},\mathbf{v}_{i})$, and consider the $N$-geodesics $\gamma_{i}:[0,\ell_{i}]\to M$ joining $p_{i}=\gamma_{i}(0)$ to $q_{i}=\gamma_{i}(\ell_{i})$, given by $\gamma_{i}(t)=\exp^{\nu}(t\mathbf{v}_{i})$. From the definition of $\rho$, we have $\gamma_{i}$ is an $N$-segment for each $i$. Since $\rho(p_{i},\mathbf{v}_{i})\in[0,\infty]$, we must have some accumulation point, say, $L_{0}\in[0,\infty]$ for the set $\{\rho(p_{i},\mathbf{v}_{i})\}$. Consider the $N$-geodesic $\gamma:[0,L_{0}]\to M$ defined as $\gamma(t)=\exp^{\nu}(t\mathbf{v})$. Passing to subsequence, we have $\mathbf{v}_{i}=\dot{\gamma}_{i}(0)\to\mathbf{v}=\dot{\gamma}(0)$, and $\ell_{i}=\rho(p_{i},\mathbf{v}_{i})\to L_{0}$. But then,
$d(N,\gamma(L_{0}))=d\left(N,\exp^{\nu}(L_{i}\mathbf{v})\right)=\lim d(N,\exp^{\nu}(\ell_{i}\mathbf{v}_{i}))=\lim d(p_{i},\gamma_{i}(\ell_{i}))=\lim\ell_{i}=L_{0},$
and thus $\gamma$ is an $N$-segment. By definition of $\rho$, we have $L_{0}\leq\rho(p,\mathbf{v})$. If $L_{0}=\infty$, then $\rho(p,\mathbf{v})=L_{0}$, and we are done.
Suppose $L_{0}<\infty$, and so for infinitely many $(p_{i},\mathbf{v}_{i})$, $q_{i}=\exp^{\nu}(\ell_{i}\mathbf{v}_{i})$ is a cut point of $N$. It follows from 4.6 that there are infinitely many $(p_{i},\mathbf{v}_{i})$ which satisfy one of the following.
  1.  (a)
    $q_{i}=\gamma_{i}(\ell_{i})=\exp^{\nu}(\ell_{i}\mathbf{v}_{i})$ is the first focal point of $N$ along $\gamma_{i}$.
  2.  (b)
    There exists some $N$-segemnt ending at $q_{i}$ distinct from $\gamma_{i}$.
If (a) holds for infinitely many $(p_{i},\mathbf{v}_{i})$, then $\ker\left(d(\exp^{\nu})|_{\ell_{i}\mathbf{v}_{i}}\right)\neq 0$. Since $\lim\ell_{i}=L_{0}$ is finite, passing to a subsequence, we may assume that there exists some $\epsilon>0$ so that $L_{0}-\epsilon<\ell_{i}<L_{0}+\epsilon$ for all $i$. Consider the subset
$\widetilde{\nu}\coloneqq\{(p,\mathbf{v})\in\nu\;|\;L_{0}-\epsilon\leq F_{p}(\mathbf{v})\leq L_{0}+\epsilon\},$
which is clearly a compact submanifold of $\hat{\nu}$ of codimension 0. Furthermore, $\ell_{i}\mathbf{v}_{i}\in\widetilde{\nu}$, and $T_{\ell_{i}\mathbf{v}_{i}}\hat{\nu}=T_{\ell_{i}\mathbf{v}_{i}}\widetilde{\nu}$. Let us now fix some auxiliary Riemannian metric on $\widetilde{\nu}$, and choose unit vectors $\mathbf{w}_{i}\in\ker\left(d(\exp^{\nu})|_{\ell_{i}\mathbf{v}_{i}}\right)$. Since $\mathbf{w}_{i}$ is in the unit sphere bundle $S(T\widetilde{\nu})$ of a compact manifold $\widetilde{\nu}$, passing to a further subsequence, we may assume $\mathbf{w}_{i}\to\mathbf{w}\in S(T\widetilde{\nu})$. Then $0=\lim d(\exp^{\nu})|_{\ell_{i}\mathbf{v}_{i}}(\mathbf{w}_{i})=d(\exp^{\nu})|_{L_{0}\mathbf{v}}(\mathbf{w})$. As $\mathbf{w}\neq 0$, $q=\gamma(T)=\exp^{\nu}(L_{0}\mathbf{v})$ is a focal point of $N$ along $\gamma$. Hence, any extension of $\gamma$ beyond $L_{0}$ fails to be distance minimizing (4.4), and we get $\rho(p,\mathbf{v})=L_{0}$.
If (b) holds for infinitely many $(p_{i},\mathbf{v}_{i})$, there exists $(p_{i}^{\prime},\mathbf{v}_{i}^{\prime})\in S(\nu)$ so that the geodesic $\eta_{i}:[0,\ell_{i}]\to M$ defined by $\eta_{i}(t)=\exp^{\nu}(t\mathbf{v}_{i}^{\prime})$ is distinct from $\gamma_{i}$, and is an $N$-segment satisfying $q_{i}=\eta_{i}(\ell_{i})$. Passing to a subsequence, we may assume $(p_{i}^{\prime},\mathbf{v}_{i}^{\prime})\to(p^{\prime},\mathbf{v}^{\prime})$. The $N$-geodesic $\gamma^{\prime}:[0,L_{0}]\to M$ defined by $\gamma^{\prime}(t)=\exp^{\nu}(t\mathbf{v}^{\prime})$ satisfies $\gamma^{\prime}(L_{0})=\exp^{\nu}(L_{0}\mathbf{v}^{\prime})=\lim\exp^{\nu}(\ell_{i}\mathbf{v}_{i}^{\prime})=\lim q_{i}=q$, and so it is an $N$-segment. If $v^{\prime}\neq v$, then we have $q\in\mathrm{Se}(N)\subset\mathrm{Cu}(N)$ (4.2), and we get $\rho(p,\mathbf{v})=L_{0}$. Next, suppose $\mathbf{v}^{\prime}=\mathbf{v}$ and so, $\gamma^{\prime}=\gamma$. But then $\exp^{\nu}$ fails to be injective in any neighborhood of $L_{0}\mathbf{v}$, since we have $(p_{i}^{\prime},\mathbf{v}_{i}^{\prime})\neq(p_{i},\mathbf{v}_{i})\Rightarrow\ell_{i}\mathbf{v}_{i}^{\prime}\neq\ell_{i}\mathbf{v}_{i}$, both of which are getting mapped to $q_{i}$ under the exponential map. In particular, $d(\exp^{\nu})|_{L_{0}\mathbf{v}}$ is non-degenerate and thus $q=\exp^{\nu}(L_{0}\mathbf{v})$ is a focal point of $N$ along $\gamma$. Then by 4.4, we get $\rho(p,\mathbf{v})=L_{0}$.
Thus, for any accumulation point $L_{0}$ of $\{\rho(p_{i},\mathbf{v}_{i})\}$ we have $L_{0}=\rho(p,\mathbf{v})$. In particular, $\rho(p_{i},\mathbf{v}_{i})\to\rho(p,\mathbf{v})$, which completes the proof. ∎
We can now prove the following.
Theorem 4.8.
[…]
Proof.
Let $q\in\mathrm{Cu}(N)\setminus\mathrm{Se}(N)$. Then, there exists a unique $N$-segment $\gamma:[0,\ell]\to M$ joining $p=\gamma(0)\in N$ to $q=\gamma(\ell)$, and furthermore by 4.6, $q$ is the first focal point of $N$ along $\gamma$. We have $q=\exp^{\nu}(\ell\mathbf{v})$, where $\mathbf{v}=\dot{\gamma}(0)\in\nu_{p}$, and $\ell=\rho(\mathbf{v})$.
[…]
$U(\mathbf{v},\delta)=\left\{\rho(\mathbf{w})\mathbf{w}\;\middle|\;\mathbf{w}\in V(\mathbf{v},\delta)\right\}.$
[…]
  1.  (a)
[…]
  2.  (b)
    For any $\delta>0$, $\exp^{\nu}|_{U(\mathbf{v},\delta)}$ is not injective.
If (b) appears, then for each $k$ we can pick $\mathbf{u}_{k}^{1}\neq\mathbf{u}_{k}^{2}\in U_{k}$, so that $q_{k}=\exp^{\nu}(\mathbf{u}_{k}^{1})=\exp^{\nu}(\mathbf{u}_{k}^{2})$. Note that $q_{k}\in\mathrm{Se}(N)$. As $\cap_{k}U_{k}=\left\{\mathbf{u}\right\}$, we have $\mathbf{u}_{k}^{1}\rightarrow u$ and $\mathbf{u}_{k}^{2}\rightarrow u$. Hence, it follows from the continuity of $\exp^{\nu}$ that $q_{k}\rightarrow q=\exp^{\nu}(u)$. As $q_{k}\in\mathrm{Se}(N)$, $q$ is in the closure of $\mathrm{Se}(N)$, completing the proof.
[…]
$V(\mathbf{v},\delta)$
$S(\nu)$
$U(\mathbf{v},\delta)$
$\mathbf{v}$
$\ell_{i}\mathbf{v}_{i}$
$\mathbf{v}_{i}$
$\exp^{\nu}$
$q$
$B(q,\varepsilon)$
$W$
$q_{i}$
$p_{i}$
$\gamma_{i}$
$\mathbf{v}_{i}=\dot{\gamma}_{i}(0)$
$\mathbf{v}=\dot{\gamma}(0)$
$\gamma$
$p$
$N$
$\Phi$
$\rho(\mathbf{v})\mathbf{v}$
Figure 3. The open sets in $\nu$ and in $M$
As $S(\nu)$ is compact, passing to a subsequence if necessary, we have $(p_{i},\mathbf{v}_{i})\rightarrow(p^{\prime},\mathbf{v}^{\prime})$. We also have,
$\ell=d(N,q)=\lim d(N,q_{i})=\lim\ell_{i}.$
Consequently, the $N$-geodesic $\eta:[0,\ell]\rightarrow M$ given by $\eta(t)=\exp^{\nu}(t\mathbf{v}^{\prime})$ is in fact an $N$-segment joining $N$ to $q$. We claim that $\gamma$ and $\eta$ are distinct. If this were not the case, we must have $(p,\mathbf{v})=(p^{\prime},\mathbf{v}^{\prime})$. But as $\mathbf{v}_{i}\rightarrow\mathbf{v}$ in $S(\nu)$, and as $V_{k_{0}}\vcentcolon=V\left(\mathbf{v},\frac{1}{k_{0}}\right)=\Phi^{-1}(U_{k_{0}})$ is an open neighborhood of $\mathbf{v}$ in $S(\nu)$, we must have $\mathbf{v}_{i}\in V_{k_{0}}$ for all $i$ sufficiently large. Consequently, $\Phi(\mathbf{v}_{i})=\ell_{i}\mathbf{v}_{i}\in U_{k_{0}}$ for all $i$ large. But that would imply $q_{i}=\exp^{\nu}(\ell_{i}\mathbf{v}_{i})\in W$, which is a contradiction. Thus, we have two distinct $N$-segments $\gamma$ and $\eta$ joining $N$ to $q$, making $q\in\mathrm{Se}(N)$, which is again a contradiction. Therefore, (a) cannot appear.
Now, suppose $N$ is closed. Then $S(\nu)$ is compact (3.3). It follows from 4.7, that we have a continuous map
$\Phi:S(\nu)$ $\rightarrow\nu$
$\mathbf{v}$ $\mapsto\rho(\mathbf{v})\mathbf{v}$
Then, $\widetilde{\mathrm{Cu}}(N)=\Phi(S(\nu))$ is a compact set, and hence, $\mathrm{Cu}(N)=\exp^{\nu}\left(\widetilde{\mathrm{Cu}}(N)\right)$ is a compact set. The proof then follows. ∎
4.3. The Topology of the Complement of Cut Locus
Let us consider the subset
[…]
and denote ^$\hat{\mathcal{I}}(N)=\mathcal{I}(N)\cap\hat{\nu}$. It is clear that $\mathcal{I}(N)\cap\widetilde{\mathrm{Cu}}(N)=\emptyset$, and $\exp^{\nu}(\mathcal{I}(N))\cap\mathrm{Cu}(N)=\emptyset$. Observe that $\mathcal{I}(N)$ is star-shapped, i.e., for any $\mathbf{v}\in\mathcal{I}(N)$ we have $t\mathbf{v}\in\mathcal{I}$ for $0\leq t\leq 1$.
Theorem 4.9.
The normal exponential map $\exp^{\nu}$ induces a
  • $C^{1}$-diffoemorphism of $\mathcal{I}(N)$ onto $M\setminus\mathrm{Cu}(N)$
  • $C^{\infty}$ diffeomorphism of ^$\hat{\mathcal{I}}(N)$ onto $M\setminus(N\cup\mathrm{Cu}(N))$.
Proof.
Let $q\in M\setminus\left(N\cup\mathrm{Cu}(N)\right)$. Then, it follows from 4.2, that there exists a unique $N$-segment $\gamma:[0,\ell]\rightarrow M$ joining $N$ to $q=\gamma(\ell)$, where $\ell=d(N,q)$. We have $\gamma(t)=\exp^{\nu}(t\mathbf{v})$, where $\mathbf{v}=\dot{\gamma}(0)\in S(\nu)$, and $\ell\mathbf{v}\in\hat{\mathcal{I}}(N)$. In particular, we have a map $\Phi:M\setminus\mathrm{Cu}(N)\rightarrow\mathcal{I}(N)$ defined as follows.
$\Phi(q)=\begin{cases}\ell\mathbf{v},&q\not\in N\cup\mathrm{Cu}(N),\text{ and }q=\exp^{\nu}(\ell\mathbf{v})\\ 0\in T_{q}N,&q\in N\end{cases}$
By construction, $\Phi$ is the required inverse map to $\exp^{\nu}$. We show that $\Phi$ is $C^{1}$ on $M\setminus\mathrm{Cu}(N)$, and is $C^{\infty}$ when restricted to $M\setminus(N\cup\mathrm{Cu}(N))$.
Let $q\not\in N\cup\mathrm{Cu}(N)$, and $\mathbf{u}=\ell\mathbf{v}=\Phi(q)\in\hat{\mathcal{I}}(N)$. It follows from 4.4 that $q$ is not a focal point of $N$ along the $N$-geodesic $\gamma(t)=\exp^{\nu}(t\mathbf{v})$. In particular, $d\left(\exp^{\nu}\right)|_{\mathbf{u}}$ is nondegenerate, and hence, it is a local diffeomorphism. By the implicit function theorem, $\Phi$ is then $C^{\infty}$ in a neighborhood of $q$. Thus, $\Phi$ is $C^{\infty}$ on $M\setminus\left(N\cup\mathrm{Cu}(N)\right)$. Since $\exp^{\nu}$ is $C^{1}$ on the 0 section of $\nu$, same argument as above gives that $\Phi$ is $C^{1}$ on $M\setminus\mathrm{Cu}(N)$. This completes the proof. ∎
As a consequence, we now have the following.
Theorem 4.10.
[…]
  1. (1)
    $M\setminus\mathrm{Cu}(N)$ strongly deformation retracts onto $N$.
  2. (2)
    $M\setminus N$ strongly deformation retracts onto $\mathrm{Cu}(N)$.
Proof.
Consider the linear homotopy $L_{s}:TM\rightarrow TM$ given by $L_{s}(\mathbf{v})=(1-s)\mathbf{v}$, for $0\leq s\leq 1$. By construction, $\mathcal{I}(N)$ is star-shaped, and thus, $L_{s}$ restricts to a homotopy of $\nu$ as well. Indeed,
$L_{0}|_{\mathcal{I}(N)}=\mathrm{Id}_{\mathcal{I}(N)},\quad L_{1}|_{\mathcal{I}(N)}=\pi|_{\mathcal{I}(N)},$
where $\pi:\nu\rightarrow N$ is the projection. Next, we define the map $h:I\times(M\setminus\mathrm{Cu}(N))\rightarrow(M\setminus\mathrm{Cu}(N))$ by
$h(s,q)\coloneqq L_{s}\circ\left(\exp^{\nu}\right)^{-1}(q).$
It follows from 4.9, that $h$ is a well-defined homotopy. Clearly, $h_{0}=h(0,\_)$ is identity on $M\setminus\mathrm{Cu}(N)$, whereas $h_{1}=h(1,\_)$ is contained in $N$. Furthermore, $h_{s}=h(s,\_)$ is identity on $N$ for all $0\leq s\leq 1$, and $h_{1}\circ\iota_{N}=\mathrm{Id}_{N}$, where $\iota_{N}:N\hookrightarrow M$ is the inclusion. Thus, $M\setminus\mathrm{Cu}(N)$ strongly deformation retracts onto $N$, which proves (1).
In order to prove (2) consider the homotopy $H:I\times(M\setminus N)\rightarrow M\setminus N$ defined as
$H(s,q)\coloneqq\begin{cases}\exp^{\nu}\left[\left(s\rho\left(\frac{\mathbf{v}}{F(\mathbf{v})}\right)+(1-s)F(\mathbf{v})\right)\frac{\mathbf{v}}{F(\mathbf{v})}\right],&q\in M\setminus\left(N\cup\mathrm{Cu}(N)\right),\\ &\mathbf{v}\coloneqq\left(\exp^{\nu}\right)^{-1}(q)\\[10.00002pt] q,&q\in\mathrm{Cu}(N).\end{cases}$
Continuity of $H$ follows from 4.7 and 4.9. Clearly, $H_{0}=H(0,\_)$ is identity on $M\setminus N$, whereas $H_{1}=H(1,\_)$ is contained in $\mathrm{Cu}(N)$. Furthermore, for all $0\leq s\leq 1$, we have $H_{s}=H(s,\_)$ is identity on $\mathrm{Cu}(N)$, and $H_{1}\circ\iota_{\mathrm{Cu}(N)}=\mathrm{Id}|_{\mathrm{Cu}(N)}$, where $\iota_{\mathrm{Cu}(N)}:\mathrm{Cu}(N)\hookrightarrow M$ is the inclusion. Thus, $M\setminus N$ strongly deformation retracts onto $\mathrm{Cu}(N)$. ∎
Observation 4.11.
Suppose $N$ is a closed submanifold of $M$. We make a few easy observations.
  1. (1)
    $M=\exp^{\nu}\left(\mathcal{I}(N)\cup\widetilde{\mathrm{Cu}}(N)\right)$.
  2. (2)
    $M$ is compact if and only if $\rho$ is finite. Indeed, if $M$ is compact, then continuity of $\rho$ (4.7) implies that $\rho$ is finite. For the converse, first, consider the unit normal disc bundle of $N$
[…]
    which is compact as $N$ is so. Since $\rho$ is finite and continuous, we have a continuous injective map $f:D(\nu)\rightarrow\nu$ defined as
    $f(\mathbf{v})=\begin{cases}\rho\left(\frac{\mathbf{v}}{F(\mathbf{v})}\right)\mathbf{v},&\mathbf{v}\neq 0\\ 0,&\mathbf{v}=0\end{cases}.$
    Note that $\operatorname{im}f=\mathcal{I}(N)\cup\mathrm{Cu}(N)$, which is compact. But then $M$ is compact by the previous observation.
  3. (3)
    If $\lambda<\infty$, then $M$ is compact, since $\rho\leq\lambda$ by 4.5. The converse may not be true, even for the Riemannian case: consider any compact Riemannian manifold $M$ with flat curvature, then any submanifold $N\subset M$ cannot have any focal locus and hence $\lambda=\infty$.
4.4. Distance Function from $N$
Theorem 4.12.
Let $N$ be a closed submanifold, and denote by $f:M\rightarrow\mathbb{R}$ the distance squared function $f(q)=d(N,q)^{2}$. Then,
  1. (1)
    $f$ is $C^{1}$ precisely on $M\setminus\mathrm{Se}(N)$, and $f$ is $C^{\infty}$ on $M\setminus(N\cup\mathrm{Cu}(N))$.
  2. (2)
    For $q\not\in\mathrm{Se}(N)$ and $X\in T_{q}M$, we have $df(X)=2\ell g_{\dot{\gamma}(\ell)}(\dot{\gamma}(\ell),X)$, where $\gamma:[0,\ell]\rightarrow M$ is the unique $N$-segment joining $N$ to $q=\gamma(\ell)$ and $\ell=d(N,q)$.
  3. (3)
    For $q\in\mathrm{Se}(N)$, $f$ fails to be differentiable precisely along those $X\in T_{q}M$ for which there exists an $N$-segment $\gamma:[0,\ell]\rightarrow M$ joining $N$ to $q$ with $\dot{\gamma}(\ell)=X$.
The proof is essentially the same as the one given in [ST16], where the authors proved that the distance function from an arbitrary closed set $N$ in a Finsler manifold is $C^{1}$ precisely in the compliment of $N\cup\mathrm{Cu}(N)$. Recall the following.
Lemma 4.13.
[ST16, Prop 2.2] Consider the following setup.
  • $\gamma:[0,\ell]\rightarrow M$ is an $N$-segment, joining $p=\gamma(0)$ to $q=\gamma(\ell)$, where $\ell=d(N,q)$.
  • $q_{i}\rightarrow q$ is a converging sequence in $M$.
  • $\gamma_{i}:[0,\ell_{i}]\rightarrow M$ is an $N$-segment, joining $p_{i}=\gamma_{i}(0)$ to $q_{i}=\gamma_{i}(\ell_{i})$, where $\ell_{i}=d(N,q_{i})$, and $\mathbf{v}_{i}=\dot{\gamma}_{i}(0)$.
  • $\sigma_{i}:[0,T_{i}]\rightarrow M$ is an $N$-segment, joining $q=\sigma_{i}(0)$ to $q_{i}=\sigma_{i}(T_{i})$, where $T_{i}=d(q,q_{i})$, and $\mathbf{w}_{i}=\dot{\sigma}_{i}(0)$.
Suppose, the limit $\mathbf{w}=\lim_{i}\mathbf{w}_{i}$ exists, and $\gamma_{i}$ converges uniformly to $\gamma$. Then, the limit
$\lim_{i\rightarrow\infty}\frac{d(N,q_{i})-d(N,q)}{d(q,q_{i})}$
exists, and equals $g_{\dot{\gamma}(\ell)}\left(\dot{\gamma}(\ell),\mathbf{w}\right)$.
Proof of 4.12.
We have from 4.9, $\exp^{\nu}$ restricts to a $C^{\infty}$-diffeomorphism ^$\hat{\mathcal{I}}(N)\rightarrow M\setminus(N\cup\mathrm{Cu}(N))$. For each $q\in M\setminus(N\cup\mathrm{Cu}(N))$, we have the unique $N$-segment $\gamma_{\mathbf{v}}$ joining $N$ to $q$, where $\mathbf{v}=(\exp^{\nu})^{-1}(q)$, and furthermore $d(N,q)=L(\gamma_{\mathbf{v}})=F(\mathbf{v})$. Thus, we have the identity
$f|_{M\setminus(N\cup\mathrm{Cu}(N))}=\left(F\circ(\exp^{\nu})^{-1}\right)^{2},$
which proves the smoothness of $f$ on the compliment of $N\cup\mathrm{Cu}(N)$.
Now, let $q\not\in\mathrm{Se}(N)$. Then, there exists a unique $N$-segment $\gamma:[0,\ell]\rightarrow M$ joining $p=\gamma(0)\in N$ to $q=\gamma(\ell)$, where $\ell=d(N,q)$ and $\mathbf{v}=\dot{\gamma}(0)\in S(\nu)$. Clearly, $f(q)=\ell^{2}=(L(\gamma))^{2}$. Let $X\in T_{q}M$, and consider a smooth curve $\sigma:(-\epsilon,\epsilon)\rightarrow M$ such that $\sigma(0)=q,\dot{\sigma}(0)=X$. Let $s_{i}\rightarrow 0$ be a sequence in the domain of $\sigma$. Suppose, $\gamma_{i}:[0,\ell_{i}]\rightarrow M$ are $N$-segments, joining $p_{i}=\gamma_{i}(0)\in N$ to $q_{i}=\sigma(s_{i})=\gamma_{i}(\ell_{i})$, where $\ell_{i}=d(N,q_{i})=L(\gamma_{i})$. As $s_{i}\rightarrow 0$, we have $q_{i}=\sigma(s_{i})\rightarrow\sigma(0)=q$, and $\ell_{i}\rightarrow\ell$. Set $\mathbf{v}_{i}=\dot{\gamma}(0)\in S(\nu)$. It follows from the compactness of $S(\nu)$ and the uniqueness of the $N$-segment $\gamma$ joining $N$ to $q$, that $(p_{i},\mathbf{v}_{i})\rightarrow(p,\mathbf{v})$, and consequently $\gamma_{i}\rightarrow\gamma$ uniformly. In particular, $\dot{\gamma}(\ell)=\lim_{i}\dot{\gamma}_{i}(\ell_{i})$. Then, by an application of the chain rule, it follows from 4.13 that the limit
$\lim_{i\rightarrow\infty}\frac{f(\sigma(s_{i}))-f(\sigma(0))}{s_{i}}=\lim_{i\rightarrow\infty}\frac{d(N,q_{i})^{2}-d(N,q)^{2}}{s_{i}}$
exists, and equals $2\ell g_{\dot{\gamma}(\ell)}(\dot{\gamma}(\ell),X)$. A careful examination of the proof of 4.13, reveals that $f$ fails to be differentiable at points of $\mathrm{Se}(N)$ precisely along those directions which is the terminal velocity of some $N$-segment. This concludes the proof. ∎
4.5. Existence of $N$-Geodesic Loops
In this section, we generalize a result from [ISS12] and [Xu15]. The proof is the same vein, with suitable modifications needed to deal with the asymmetric nature of Finsler distance.
Theorem 4.14.
[…]
  • There are points in $\mathrm{Cu}(N)$ which are also focal points of $N$.
  • For each $q\in M$, there exists at least two $N$-geodesics joining $N$ to $q$, one of which crosses the cut locus $\mathrm{Cu}(N)$. For $q\in N$, we allow the constant curve at $q$ to be considered as an $N$-geodesic.
For any $q\in M$, consider the function
$\mathcal{M}=\mathcal{M}_{q}:M$ [0,$\rightarrow$[0$\infty$)→ [ 0 , ∞ )
$x$ $\mapsto d(N,x)+d(x,q)$
Since $N$ is closed, we have $\mathrm{Cu}(N)$ is closed (4.8), and consequently, there exists some $x_{0}\in\mathrm{Cu}(N)$ so that $\mathcal{M}|_{\mathrm{Cu}(N)}$ attains a global minima at $x_{0}$, i.e.,
$\mathcal{M}(x_{0})=\min_{x\in\mathrm{Cu}(N)}\mathcal{M}(x).$
We have the following lemma, which is quite interesting in itself.
Lemma 4.15.
With the above notation, further assume that $x_{0}$ is not a focal point of $N$. We then have the following dichotomy.
  1. (1)
    $q\not\in\mathrm{Cu}(x_{0})$. Then, there exists exactly two $N$-segments, say, $\gamma_{1},\gamma_{2}$, joining $N$ to $q$, and a unique minimizer, say $\eta$, joining $x_{0}$ to $q$. Furthermore, exactly one of $\gamma_{1}\cup\eta$ and $\gamma_{2}\cup\eta$ is an $N$-geodesic joining $N$ to $q$, and intersecting $\mathrm{Cu}(N)$ at $x_{0}$.
  2. (2)
    $q\in\mathrm{Cu}(x_{0})$. Then, there exists exactly two $N$-segments, say, $\gamma_{1},\gamma_{2}$, joining $N$ to $x_{0}$, and at most two minimizers, say, $\eta_{1},\eta_{2}$, joining $x_{0}$ to $q$. Furthermore, if $\eta_{1},\eta_{2}$ are distinct, then exactly two of the four curves $\left\{\gamma_{i}\cup\eta_{j}\;\middle|\;i,j=1,2\right\}$ are $N$-geodesics joining $N$ to $q$ and crossing $\mathrm{Cu}(N)$. The other two are broken at $x_{0}$. If $\eta_{1}=\eta_{2}=\eta$, then exactly one of $\left\{\gamma_{1}\cup\eta,\gamma_{2}\cup\eta\right\}$ is an $N$-geodesic joining $N$ to $q$ crossing $\mathrm{Cu}(N)$ at $x_{0}$.
  3. (2’)
    For $q\in\mathrm{Cu}(x_{0})$, if we furthermore assume that $F$ is a reversible Finsler metric, then exactly two minimizers exist from $x_{0}$ to $q$, and the rest of (2) holds.
Proof.
Case (1): Suppose $q\not\in\mathrm{Cu}(x_{0})$. Then, it follows from 4.3 that there exists a unique minimizer $\eta:[0,a]\rightarrow M$ joining $x_{0}$ to $q$, where $a=d(x_{0},q)$. If possible, assume that we have at least three $N$-segments joining $N$ to $x_{0}$, say, $\gamma_{1},\gamma_{2},\gamma_{3}:[0,\ell]\rightarrow M$, where $\ell=d(N,x_{0})$ (Figure 4). Since they are distinct, their terminal velocities $\dot{\gamma}_{i}(\ell)$ are distinct as well (2.10). Consequently, at most one of $\gamma_{i}\cup\eta$ can be smooth at $x_{0}$. Without loss of generality, assume that both $\gamma_{1}\cup\eta$ and $\gamma_{2}\cup\eta$ are broken at $x_{0}$.
$x_{0}$
$\mathrm{Cu}(N)$
$q\notin\mathrm{Cu}(x_{0})$
$\gamma_{1}$
$\gamma_{3}$
$\gamma_{2}$
$N$
$T$
Figure 4. Case (1)
Since $\gamma_{1}$ is in particular a minimizer, $\gamma_{1}|_{[t,\ell]}$ is a minimizer for any $0<t<\ell$ (2.13). Then, for $y=\gamma_{1}(t)$, it follows from the brokenness of $\gamma_{1}\cup\eta$ that
$d(y,q)<d(y,x_{0})+d(x_{0},q).$
Hence,
$\mathcal{M}(y)=d(N,y)+d(y,q)<d(N,y)+d(y,x_{0})+d(x_{0},q)=d(N,x_{0})+d(x_{0},q)=\mathcal{M}(x_{0}).$
[…]
$x_{0}$
$\mathrm{Cu}(N)$
$q\notin\mathrm{Cu}(x_{0})$
$\gamma_{1}$
$\gamma_{3}$
$\gamma_{2}$
$N$
$T$
$y$
$\zeta$
$\zeta$
$x_{0}^{\prime}$
Figure 5. Case (1) : Minimizer $\zeta$ joining $y$ to $q$, crossing $\mathrm{Cu}(N)$
We have
$\mathcal{M}(x_{0})\leq\mathcal{M}(x_{0}^{\prime})$ $=d(N,x_{0}^{\prime})+d(x_{0}^{\prime},q)$
$\leq d(N,y)+\underbrace{d(y,x_{0}^{\prime})+d(x_{0}^{\prime},q)}_{\text{$y,x_{0}^{\prime},q$ are on the minimizer $\zeta$}}$
$=d(N,y)+d(y,q)=\mathcal{M}(y)<\mathcal{M}(x_{0}),$
which is a contradiction. Thus, $\zeta$ does not intersect $\mathrm{Cu}(N)$. Similarly, for any $y=\gamma_{2}(t)$, we have that any minimizer joining $y$ to $q$ does not intersect $\mathrm{Cu}(N)$ (Figure 6).
$x_{0}$
$\mathrm{Cu}(N)$
$q\notin\mathrm{Cu}(x_{0})$
$\gamma_{1}$
$\gamma_{3}$
$\gamma_{2}$
$N$
$T$
$y_{1j}$
$\zeta_{1j}$
Figure 6. Case (1) : Minimizer between $y_{ij}$ and $q$
Pick some sequence $t_{j}\rightarrow\ell$ and set $y_{ij}=\gamma_{i}(t_{j}),i=1,2$. Then, $y_{ij}\rightarrow\gamma_{i}(\ell)=x_{0}$ for $i=1,2$. Consider minimizers $\zeta_{ij}$ joining $y_{ij}$ to $q$. Since $\zeta_{ij}$ does not intersect $\mathrm{Cu}(N)$, it follows from 4.9 that there exists curves $\tilde{\zeta}_{ij}=\left(\exp^{\nu}\right)^{-1}(\zeta_{ij})$ in $\mathcal{I}(N)$ joining $\tilde{y}_{ij}=\left(\exp^{\nu}\right)^{-1}(y_{ij})$ to $\tilde{q}=\left(\exp^{\nu}\right)^{-1}(q)$. As $x_{0}$ is not a focal point of $N$, it follows that $\exp^{\nu}$ is a local diffeomorphism near $\ell\dot{\gamma}_{i}(0)\in\left(\exp^{\nu}\right)^{-1}(x_{0})$ for $i=1,2$. In particular, passing to a subsequence if necessary, we have $\tilde{y}_{ij}\rightarrow\ell\dot{\gamma}_{i}(0)$ for $i=1,2$. By an Arzelà–Ascoli type argument ([Bus55, Theorem 5.16]), we then have (possibly a subsequence of) $\tilde{\zeta}_{ij}$ converges uniformly to some $\tilde{\zeta}_{i}$ in $\mathcal{I}(N)\cup\widetilde{\mathrm{Cu}}(N)$ for $i=1,2$, which must join $\ell\dot{\gamma}_{i}(0)$ to $\tilde{q}$, respectively. Suppose, $\zeta_{i}=\exp^{\nu}(\tilde{\zeta}_{i})$, which then joins $x_{0}$ to $q$ for $i=1,2$. Note that
$L(\zeta_{i})=L\left(\exp^{\nu}\left(\tilde{\zeta}_{i}\right)\right)=L\left(\lim_{j}\exp^{\nu}(\tilde{\zeta}_{ij})\right)=L\left(\lim_{j}\zeta_{ij}\right)=\lim_{j}d(y_{ij},q)=d(x_{0},q),$
and hence, $\zeta_{i}$ is a minimizer joining $x_{0}$ to $q$. From uniqueness, we have $\zeta_{1}=\eta=\zeta_{2}$, up to a reparametrization. We now show that this leads to a contradiction.
Consider a reparametrization of $\tilde{\zeta}_{i}$ given as $\tilde{\eta}_{i}:[0,a]\rightarrow\nu$, where $a=d(x_{0},q)$, so that $\exp^{\nu}(\tilde{\eta}_{i}(t))=\eta(t)$ for $0\leq t\leq a$ and $i=1,2$. Set
$s_{0}:=\inf_{0\leq s\leq a}\left\{s\;\middle|\;\tilde{\eta}_{1}|_{[s,a]}=\tilde{\eta}_{2}|_{[}s,a]\right\}.$
[…]
Thus, we conclude that there exists exactly two $N$-segments joining $N$ to $x_{0}$. Suppose, $\gamma_{1},\gamma_{2}:[0,\ell]\rightarrow M$ are the $N$-segments. We must have exactly one of $\gamma_{i}\cup\eta$ is smooth, since otherwise by the above argument we will again reach a contradiction. Without loss of generality, assume that $\gamma_{1}\cup\eta$ is smooth at $x_{0}=\gamma_{1}(\ell)=\eta(0)$, i.e., $\dot{\gamma}_{1}(\ell)=\dot{\eta}(0)$. But then it follows from the uniqueness of the solution to the geodesic equation that $\gamma_{1}\cup\eta$ is an $N$-geodesic. This proves (1).
Case (2): In order to prove (2), assume that $q\in\mathrm{Cu}(x_{0})$. Say, $\eta:[0,a]\rightarrow M$ is a minimizer joining $x_{0}$ to $q$, where $a=d(x_{0},q)$. We can choose some $0<a_{0}<a$, so that $q_{0}=\eta(a_{0})\not\in\mathrm{Cu}(x_{0})$. Consider the function $\mathcal{M}_{0}=\mathcal{M}_{q_{0}}$ defined as $\mathcal{M}_{0}(x)\vcentcolon=d(N,x)+d(x,q_{0})$ for $x\in M$. We show that $\mathcal{M}_{0}(x_{0})=\min_{x\in\mathrm{Cu}(N)}\mathcal{M}_{0}(x)$. If not, suppose there exists some $x_{1}\in\mathrm{Cu}(N)$ such that $\mathcal{M}_{0}(x_{1})<\mathcal{M}_{0}(x_{0})$. Note that
$\mathcal{M}(x_{1})$ $=d(N,x_{1})+d(x_{1},q)$
$\leq d(N,x_{1})+d(x_{1},q_{0})+d(q_{0},q)$
$=\mathcal{M}_{0}(x_{1})+d(q_{0},q)$
$<\mathcal{M}_{0}(x_{0})+d(q_{0},q)$
$=d(N,x_{0})+\underbrace{d(x_{0},q_{0})+d(q_{0},q)}_{\text{$x_{0},q_{0},q$ are on the minimizer $\eta$}}$
$=d(N,x_{0})+d(x_{0},q)$
$=\mathcal{M}(x_{0}),$
which is a contradiction as $\mathcal{M}$ achieves a global minima value on $\mathrm{Cu}(N)$ at $x_{0}$. Hence, $\mathcal{M}_{0}$ also achieves a global minima on $\mathrm{Cu}(N)$ at $x_{0}$. But now we can argue as in (1), for $q_{0}$ and $\mathcal{M}_{0}$ instead of $q$ and $\mathcal{M}$ respectively. In particular, we have exactly two $N$-segments joining $N$ to $x_{0}$, say, $\gamma_{1},\gamma_{2}$.
Now, we have a minimizer joining $x_{0}$ to $q$. Suppose we have at least three distinct minimizers from $x_{0}$ to $q$. Then, we have at least one geodesic, say, $\eta$ from $x_{0}$ to $q$ such that both $\gamma_{1}\cup\eta$ and $\gamma_{2}\cup\eta$ are broken at $x_{0}$. But then picking some $q_{0}\neq x_{0},q$ on $\eta$, and then arguing as in (1) for $\mathcal{M}_{0}$ and $q_{0}$ as above, we again reach a contradiction. Hence, there exists at most two minimizers, say, $\eta_{1},\eta_{2}$, from $x_{0}$ to $q$. If $\eta_{1},\eta_{2}$ are distinct, then same argument gives that two (and hence, exactly two) of the curves $\left\{\gamma_{i}\cup\eta_{j},\;i,j=1,2\right\}$ are smooth, which in turn gives two $N$-geodesics, both crossing $\mathrm{Cu}(N)$ at $x_{0}$. The other two curves are broken at $x_{0}$. If $\eta_{1}=\eta_{2}=\eta$, then we are in the same situation as (1), and the conclusion follows. This completes the proof of (2).
Case (2’): Lastly, let us assume that $F$ is a reversible Finsler metric so that the distance is now symmetric. In particular, $q\in\mathrm{Cu}(x_{0})\Rightarrow x_{0}\in\mathrm{Cu}(q)$. If possible, let $\eta$ be the unique minimizer joining $x_{0}$ to $q$, and without loss of generality, $\gamma_{1}\cup\eta$ be smooth at $x_{0}$. Then, for any point $y\neq x_{0}$ on $\gamma_{1}$, the minimizer joining $y$ to $q$ cannot be contained in $\gamma_{1}\cup\eta$, otherwise this would extend the reversed minimizer $\bar{\eta}$, contradicting $x_{0}\in\mathrm{Cu}(N)$. Consequently, $d(y,q)<d(y,x_{0})+d(x_{0},q)$. Since $\gamma_{2}\cup\eta$ is broken at $x_{0}$, for any $y\neq x_{0}$ on $\gamma_{2}$, we again have $d(y,q)<d(y,x_{0})+d(x_{0},q)$. But now we are in the same situation as in case (1), and arguing similarly we reach a contradiction. Thus, $\eta$ cannot be the unique minimizer joining $x_{0}$ to $q$. Arguing as in (2), we conclude that there are exactly two such minimizers, and this finishes the proof. ∎
Proof of 4.14.
Suppose $\mathrm{Cu}(N)$ has no focal points of $N$ so that we can appeal to 4.15. If $q\in\mathrm{Cu}(N)$, then by 4.6, there exists at least two $N$-geodesics (in fact, $N$-segments) joining $N$ to $q$. Suppose $q\not\in\mathrm{Cu}(N)$, so that we have a (unique) $N$-segment joining $N$ to $q$, which clearly does not intersect $\mathrm{Cu}(N)$. Note that if $q\in N$, we consider the constant curve at $q$ as an $N$-segment. Also, from 4.15 we get at least one $N$-geodesic joining $N$ to $q$, which crosses $\mathrm{Cu}(N)$, and thus distinct from the $N$-segment we already have. This concludes the claim. ∎
Following [Xu15], we have the following generalization.
Lemma 4.16.
Suppose $x_{0}$ is a local minima of the function $\mathcal{M}|_{\mathrm{Cu}(N)}$. If $x_{0}$ is not a focal point of $N$, then there exists at least two $N$-geodesics joining $N$ to $q$, one of which crosses $\mathrm{Cu}(N)$ at $x_{0}$.
Proof.
Let $\eta:[0,a]\rightarrow M$ be a minimizer joining $x_{0}$ to $q$, where $a=d(x_{0},q)$. For any $z=\eta(t)$, we have
$\mathcal{M}_{z}(x_{0})=d(N,x_{0})+d(x_{0},z)=d(N,x_{0})+d(x_{0},q)-d(z,q)=\mathcal{M}_{q}(x_{0})-d(z,q).$
Then, for $x\in\mathrm{Cu}(N)$ sufficiently close to $x_{0}$ we see,
$\mathcal{M}_{z}(x)$ $=d(N,x)+d(x,z)$
$\geq d(N,x)+d(x,q)-d(z,q)$
$=\mathcal{M}_{q}(x)-d(z,q)$
$\geq\mathcal{M}_{q}(x_{0})-d(z,q)$
$=\mathcal{M}_{z}(x).$
In particular, $x_{0}$ is again a local minima of $\mathcal{M}_{z}|_{\mathrm{Cu}(N)}$. We show that for some $z=\eta(t)$, $\mathcal{M}_{z}|_{\mathrm{Cu}(N)}$ attains a global minima at $x_{0}$. In fact, it suffices to show that $\mathcal{M}_{z}|_{\mathrm{Cu}(N)}$ attains its global minima sufficiently close to $x_{0}$ for some $z$, whence, $x_{0}$ being a local minima will imply it is a global minima as well.
Fix some $t_{j}\rightarrow 0$ and denote $q_{j}=\eta(t_{j})$, so that $q_{j}\rightarrow x_{0}$. If possible, suppose, for some open neighborhood $U\subset\mathrm{Cu}(N)$ of $x_{0}$, we have that $\mathcal{M}_{q_{j}}|_{\mathrm{Cu}(N)}$ attains a global minima, say, $z_{j}\in\mathrm{Cu}(N)\setminus U$. As $\mathrm{Cu}(N)$ is compact, passing to a subsequence, we assume that $z_{j}\rightarrow z_{0}\in\mathrm{Cu}(N)\setminus U$. Now, for any $y\in\mathrm{Cu}(N)$ we have
$d(N,z_{j})+d(z_{j},q_{j})=\mathcal{M}_{q_{j}}(z_{j})\leq\mathcal{M}_{q_{j}}(z_{j})\leq\mathcal{M}_{q_{j}}(y)=d(N,q_{j})+d(q_{j},y).$
Taking $j\rightarrow\infty$ we have,
$d(N,z_{0})+d(z_{0},x_{0})\leq d(N,x_{0})+d(x_{0},y)\Rightarrow\mathcal{M}_{x_{0}}(z_{0})\leq\mathcal{M}_{x_{0}}(y),\quad\forall y\in\mathrm{Cu}(N).$
In particular, taking $y=x_{0}$ we get
$\mathcal{M}_{x_{0}}(z_{0})\leq\mathcal{M}_{x_{0}}(x_{0})$
$\Rightarrow$ $d(N,z_{0})+d(z_{0},x_{0})\leq d(N,x_{0})+0\leq d(N,z_{0})+d(z_{0},x_{0})$
$\Rightarrow$ $d(N,x_{0})=d(N,z_{0})+d(z_{0},x_{0}).$
In particular, $z_{0}$ must be on an $N$-segment joining $N$ to $x_{0}$. But $z_{0}\in\mathrm{Cu}(N)$, and hence any $N$-segment connecting $N$ to $z_{0}$ cannot be an $N$-segment after $z_{0}$. In particular $z_{0}=x_{0}$, which contradicts $z_{0}\not\in U$. Thus, we must have for some $q_{j_{0}}=\eta(t_{j_{0}})$, $\mathcal{M}_{q_{j_{0}}}|_{\mathrm{Cu}(N)}$ attains a global minima at $x_{0}$.
Now, we can apply 4.15 for $\mathcal{M}_{q_{j_{0}}}$, and conclude the proof. ∎
Let us now state the following result, originally due to Klingenberg in the Riemannian case for a point, and later generalized in [Xu15].
Theorem 4.17.
[…]
  • $x_{0}$ is a focal point of $N$.
  • There exists a unit-speed $N$-geodesic loop, crossing the $\mathrm{Cu}(N)$ at $x_{0}$, at the midpoint
Proof.
Suppose $x_{0}$ is not a focal point of $N$. Pick some $q\in N$ so that $d(q,x_{0})=d(N,x_{0})$. Since $F$ is reversible, we have
$\mathcal{M}_{q}(x_{0})=d(N,x_{0})+d(x_{0},q)=d(N,x_{0})+d(q,x_{0})=2d(N,x_{0}).$
Then, for $x\in\mathrm{Cu}(N)$ sufficiently close to $x_{0}$, we have from 2.8 that
$\mathcal{M}_{q}(x)=d(N,x)+d(x,q)$ $=d(N,x)+d(q,x)$
$\geq d(N,x)+d(N,x)-\underbrace{d(N,q)}_{0}=2d(N,x)$
$\geq 2d(N,x_{0})=\mathcal{M}_{q}(x_{0}),$
i.e., $x_{0}$ is a local minima of $\mathcal{M}_{q}|_{\mathrm{Cu}(N)}$. By 4.16, we then have an $N$-geodesic joining $N$ to $q\in N$ (i.e., an $N$-geodesic loop), which intersects $\mathrm{Cu}(N)$ at $x_{0}$ at time $d(N,x_{0})$, which is the midpoint. ∎
Acknowledgments
The authors would like to express their gratitude to S. Basu and J. Itoh for many fruitful discussions and suggestions in the preparation of this article. They would also like to thank ICTS for the wonderful hosting and a research friendly environment during the workshop Dualities in Topology and Geometry, where the majority of this work was done. The first author was supported by the NBHM grant no. 0204/1(5)/2022/R&D-II/5649 and the second author was supported by ICTS postdoctoral fellowship.